9. Determine and Draw the quantitative shear, bending moment diagrams and qualitative deflected curve for the following questions by using the force methodB С 86 kN/m 5 m 3 m A OOO FM
Answer #1

Solution Apply minimum potential energy (force method) to determine the Redundant Reaction (RA). 4. He B 5m Mc 86 KW m 3m TaxNow from eq ③ we get As - 5 387 +R9:2).x dx 0 = 0 + jenih 2E I RA =0 - 1 /-387.2² whow JEI 0 387 * 5 2 2 387 X 5 => 3 Answerfor member BC :- (03x 5) My = - 338 +84 (*) 337 + 116.1 (x) SFx = FRA = 116.1 At 1=0 => SFB a 116.1 KN / MB = -387 At 20 = 5